### v1

# PEFTGuard: Detecting Backdoor Attacks Against Parameter-Efficient Fine-Tuning

Abstract
Fine-tuning is an essential process to improve the performance of Large Language Models (LLMs) in specific domains, with Parameter-Efficient Fine-Tuning (PEFT) gaining popularity due to its capacity to reduce computational demands through the integration of low-rank adapters. These lightweight adapters, such as LoRA, can be shared and utilized on open-source platforms. However, adversaries could exploit this mechanism to inject backdoors into these adapters, resulting in malicious behaviors like incorrect or harmful outputs, which pose serious security risks to the community. Unfortunately, few of the current efforts concentrate on analyzing the backdoor patterns or detecting the backdoors in the adapters.
To fill this gap, we first construct (and will release) PADBench, a comprehensive benchmark that contains $13,300$ benign and backdoored adapters fine-tuned with various datasets, attack strategies, PEFT methods, and LLMs. Moreover, we propose $\mathsf{PEFTGuard}$, the first backdoor detection framework against PEFT-based adapters. Extensive evaluation upon PADBench shows that $\mathsf{PEFTGuard}$ outperforms existing detection methods, achieving nearly perfect detection accuracy ($100\%$) in most cases. Notably, $\mathsf{PEFTGuard}$ exhibits zero-shot transferability on three aspects, including different attacks, PEFT methods, and adapter ranks. In addition, we consider various adaptive attacks to demonstrate the high robustness of $\mathsf{PEFTGuard}$. We further explore several possible backdoor mitigation defenses, finding fine-mixing to be the most effective method. We envision our benchmark and method can shed light on future LLM backdoor detection research.
Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) by demonstrating remarkable capabilities across a diverse range of tasks such as text generation [4, 8], code generation [6], translation [59], and mathematical reasoning [78].
Although LLMs possess impressive in-context learning capabilities [4], fine-tuning is vital to enhance the model’s performance in understanding specific domain knowledge or better aligning with human preferences. Given the substantial number of parameters in LLMs, the widely adopted PEFT technologies, such as LoRA [23] and DoRA [37], significantly improve the adaptability of LLMs to these tasks by adjusting a limited number of parameters, thus reducing resource consumption [20]. Besides, the diverse downstream capabilities of LLMs can be enhanced by directly applying various efficient tuning adapters [77, 24, 72].
Owing to the effective capabilities and straightforward usability of the adapters, users are willing to share their well-trained adapters on open-source platforms, facilitating broader community utilization. By January 2024, the number of adapters in huggingface [^1] https://huggingface.co/models. has exceeded $10,000$, with downloads reaching over $100,000$ [12]. However, the adaptability of LLMs also introduces significant challenges and vulnerabilities. One of the critical security concerns associated with LLMs is their susceptibility to backdoor attacks [71, 66, 33, 67]. Even more concerning is that the shareable and plug-and-play characteristics of the PEFT-based adapters allow adversaries to maliciously propagate the backdoored adapters [36]. Consequently, when users incorporate these backdoored adapters into the benign LLMs, the backdoors are also integrated, leading to malicious behaviors, such as incorrect or toxic responses.
Currently, backdoor defense strategies of NLP primarily focus on detection methods [65, 2, 40, 38] and mitigation methods [51, 75, 35]. The detection methods can be classified into trigger generation [2], attention analysis [40], trigger inversion [38], and meta neural analysis [65]. These methods are primarily designed for NLP tasks involving logits-based classification, such as BERT [11], which uses the [CLS] token embedding for classification. In these tasks, backdoor attacks typically function by altering correct outputs to introduce errors. However, their effectiveness in generation tasks lacks comprehensive assessment, as the variable, context-dependent outputs, and representation vectors [60] may make consistent backdoor triggers more challenging to identify.
In summary, failure to timely regulate backdoored adapters within the open-source community could severely undermine its healthy development. Considering the characteristics of efficient tuning adapters that can propagate backdoors and the limitations of current backdoor detection methods when applied to NLP generation tasks, there is a critical need to develop a specialized backdoor detection approach tailored to PEFT-based fine-tuning in LLMs.
Our Work
Backdoor Vulnerabilities in PEFT-based Adapters. Due to the lack of awareness in the current community about the dangers that PEFT-based adapters can be used to propagate backdoor attacks, we conduct the first comprehensive analysis of the security vulnerabilities of PEFT-based adapters across different attacking scenarios. Specifically, we consider a variety of generation task datasets, including classification analysis (IMDB [41] and AG News [74]), question answering (SQuAD [53]), and instruction-following (toxic-backdoors-alpaca [15] and toxic-backdoors-hard [16]). In addition, for comprehensive evaluation, we consider different textual backdoor attacks (InsertSent [9], RIPPLES [33], Syntactic [49], and StyleBkd [48]), various PEFT methods (LoRA [23], QLoRA [10], DoRA [37], LoRA+ [21], and AdaLoRA [73]). Furthermore, we consider different types of base LLMs [60, 43, 3, 14, 39] and different training settings of PEFT, including the ranks of adapters and the target projection matrices. Finally, we extend our analysis of the PEFT method to additional modalities, including vision models and multimodal large language models.
Backdoored Adapter Detection Benchmark. To fill the gap of lacking a systematic benchmark in the domain of backdoor detection for PEFT-tuned LLMs, we construct a comprehensive dataset namely PADBench. [^2] We will open-source this dataset to advance research in this field. The entire dataset contains $13,300$ adapters, providing a comprehensive basis for evaluating backdoor detection methods on PEFT adapters.
Backdoor Detection Framework. In order to efficiently identify the backdoored adapters, we propose $\mathsf{PEFTGuard}$, the first framework specifically designed to detect backdoors within the PEFT-based adapters of LLMs. For instance, $\mathsf{PEFTGuard}$ transforms the adapters’ weights through Feature Transformation (refer to Section 4.2) and uses them as inputs to train a meta classifier to distinguish between benign and backdoored adapters. Notably, the advantages of $\mathsf{PEFTGuard}$ include not requiring additional input data or merging adapters back into the original LLMs for inference. Meanwhile, $\mathsf{PEFTGuard}$ can achieve high detection performance in a zero-shot manner.
High Detection Performance. Through comprehensive experiments on PADBench, we demonstrate that $\mathsf{PEFTGuard}$ surpasses the current State-Of-The-Art (SOTA) detection methods, achieving $99\%$ detection accuracy and $1.0$ AUC in classification tasks, and $100\%$ detection accuracy and $1.0$ AUC in generation tasks, respectively. Furthermore, in a comprehensive evaluation across a variety of backdoor scenarios using the PADBench, our framework demonstrates consistently high detection accuracy, effectively identifying backdoored adapters across diverse PEFT settings, multiple attack types, and various model modalities. Notably, $\mathsf{PEFTGuard}$ exhibits zero-shot transferability without the need for fine-tuning the detection model, effectively detecting adapters from unknown attacks.
Robustness of PEFTGuard. We further demonstrate the robustness of $\mathsf{PEFTGuard}$ against five adaptive attacks, including Gaussian Noise, FGSM [17], I-FGSM [32], PGD [42], and C&W [5]. Considering that our detection framework can be seamlessly integrated with backdoor mitigation strategies, we explore various potential mitigation methods, including Supervised Fine-Tuning (SFT), DPO [51], and Fine-mixing [75], to eliminate backdoors injected by PEFT methods, with Fine-mixing proving most effective. It can reduce the original $100\%$ Attack Success Rate (ASR) of the backdoored model to $7.2\%$ while maintaining the model performance (clean accuracy is $96.12\%$).
Our Contributions. We make the following contributions:
  • We conduct the first in-depth and comprehensive analysis, revealing the security vulnerabilities of injecting backdoors into models across different modalities using PEFT-based adapters in diverse tasks.
  • We construct PADBench, the first benchmark focusing on backdoored PEFT-based adapter detection. PADBench contains a total of $13,300$ adapters generated from multiple attack scenarios.
  • We propose $\mathsf{PEFTGuard}$, a powerful backdoor detection framework against PEFT-based adapters. Notably, $\mathsf{PEFTGuard}$ introduces a meta classifier to effectively detect backdoored adapters in a zero-shot manner.
  • Benefit from PADBench, comprehensive evaluation results demonstrate that $\mathsf{PEFTGuard}$ achieves superior detection performance, strong transferability, and high robustness.
Preliminary
LLMs
Large language models typically refer to Transformer-based [61] Pre-trained Language Models (PLMs) that contain billions (B) of parameters, such as GPT-3 (175B parameters) [4] and Llama family (more than 7B parameters) [60]. These models can be categorized into three types based on their structures:
  1. 1.
    Encoder-only PLMs only include the encoder network of Transformers, originating from BERT [11] and later evolving into models with more parameters like Roberta [39] and Deberta [22]. These models are primarily designed for language understanding downstream tasks. During the pre-training process, encoder-only PLMs leverage Masked Language Modeling (MLM) paradigm, where a certain percentage of tokens in the training samples are randomly replaced with a special symbol [MASK]. For instance, given a training sequence, the model should learn to predict the masked token by using the following cross-entropy loss:
    $\mathcal{L}_{enc}=-\sum_{i=1}^{M}\log P(x^{\text{mask}}_{i}|x_{\text{context}}),$ (1)
    where $x^{\text{mask}}_{i}$ represents the masked token and $x_{\text{context}}$ represents its context. $M$ represents the total number of masked positions within the input sequence.
  2. 2.
    Decoder-only PLMs are widely used by the most popular LLMs, including ChatGPT [46], GPT-4 [47], and Llama-3 [43], because their pre-training methods are suitable for text generation tasks. For instance, the pre-training task of decoder-only PLMs is autoregressive language modeling, using a cross-entropy loss defined as:
    $\mathcal{L}_{dec}=-\sum_{t=1}^{N}\log P(x_{t}|x_{1},x_{2},...,x_{t-1}),$ (2)
    where $P$ refers to the probability of predicting the current token $x_{t}$ given all previous tokens $x_{1},...,x_{t-1}$. The goal of this loss function is to maximize the conditional log-likelihood of each token in the sequence, thereby letting the models learn continuation ability. $N$ represents the total number of words or tokens in the sequence.
  3. 3.
    Encoder-Decoder PLMs can handle both language understanding and generation tasks since all NLP tasks can be viewed as sequence-to-sequence generation tasks [52]. Representative Encoder-Decoder PLMs include T5 [52], BART [34], and ChatGLM [14]. Their pre-training task is sequence-to-sequence modeling whose loss function can be defined as:
    $\mathcal{L}_{enc-dec}=-\sum_{j=1}^{L}\log P(y_{j}|y_{1},...,y_{j-1};\mathbf{X}),$ (3)
    where $\mathbf{X}$ is the input sequence and $y_{j}$ is the word in the target sequence. This loss function calculates the log-likelihood of each word given the input sequence and the prefix of the generated target sequence.
PEFT Methods
Overview. Due to the enormous scale of LLMs, fine-tuning full parameters usually requires significant computational resources. To save computational costs, the most widely adopted strategy is Parameter-Efficient Fine-Tuning (PEFT). In this paper, we focus on reparameterized PEFT methods, especially on LoRA[23], QLoRA [10], LoRA+ [21], AdaLoRA [73], and DoRA [37]. As illustrated in Figure 1, these PEFT methods achieve fine-tuning efficiency by introducing an additional low-rank adapter (denoted as $\Delta$) while keeping the original model frozen. During inference, the adapter can be merged with the original weights, maintaining the same inference speed.
Formulation of Adapter. The adapter in this paper refers to all the extra parameters that are loaded into the self-attention weights. Formally, assume that an LLM contains $L$ self-attention layers, so the adapter $\Delta$ stands for a collection of additional parameters applied to each layer:
$\Delta:=\{\Delta^{(1)},...,\Delta^{(l)},...,\Delta^{(L)}\}.$
Meanwhile, each self-attention layer involves four key weight matrices: query ($W_{q}$), key ($W_{k}$), value ($W_{v}$), and output ($W_{o}$). For the training process, their tuned additional parameters are denoted as $\Delta_{q}$, $\Delta_{k}$, $\Delta_{v}$, and $\Delta_{o}$, corresponding to the original model parameters $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$, respectively. Thus, $\Delta^{(l)}$ can be formulated as:
$\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\},~{}l=1,...,L.$
In the following parts, we will introduce how to generate a unit adapter (e.g., $\Delta^{(l)}_{q}$) through different PEFT methods. For the sake of brevity, we uniformly use $\Delta$ to denote a unit adapter.
Figure 1: Illustration of the reparameterization PEFT algorithm.
LoRA [23]. Given a layer weight $W_{0}\in\mathbb{R}^{d\times k}$, LoRA decomposes it into two low-rank matrices, $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, which will together form the layer-specific adapter $\Delta=BA$. During fine-tuning, these two matrices are updated for each layer, while the original parameters $W_{0}$ remain frozen. For the input $\mathbf{x}\in\mathbb{R}^{k\times 1}$, the computation process of the forward pass at layer $i$ is
$\mathbf{y}=W_{0}\mathbf{x}+\Delta\mathbf{x}=W_{0}\mathbf{x}+BA\mathbf{x}.$
QLoRA [10]. QLoRA introduces several new techniques, including 4-bit NormalFloat, double quantization, and paged optimizers, which propagate 4-bit quantized pre-trained language models backward into LoRA, significantly reducing memory usage. In this context, the adapter $\Delta$ is implemented in LoRA’s low-rank form as defined above. The training process can be defined as follows:
$\begin{split}\mathbf{Y}^{\mathrm{BF}16}=&\ \mathbf{X}^{\mathrm{BF}16}doubleDequant(c_{1}^{\mathrm{FP}32},c_{2}^{\mathrm{k}-\mathrm{bit}},\mathbf{W}^{NF4})\ \\ &+\mathbf{X}^{\mathrm{BF}16}\Delta^{BF16},\end{split}$
where $\Delta^{BF16}=B^{BF16}A^{BF16}$ represents the low-rank matrices in BF16 (Bfloat16) format. Here, $doubleDequant(\cdot)$ represents a double dequantization process that first dequantizes $c_{1}^{\mathrm{FP}32}$ and $c_{2}^{\mathrm{k}\text{-bit}}$ to an intermediate representation, which is then further dequantized with $\mathbf{W}^{4\mathrm{bit}}$ to obtain the final BF16 matrix $\mathbf{W}^{\mathrm{BF}16}$. $\mathbf{X}^{\mathrm{BF}16}$ and $\mathbf{Y}^{\mathrm{BF}16}$ represent the input and output in BF16 format, respectively.
LoRA+ [21]. LoRA+ suggests setting different learning rates for the two matrices $B$ and $A$ that comprise $\Delta$, denoted as $\Delta=BA$. Specifically, $\eta_{B}=\lambda\eta_{A}$ ($\eta$ represents the learning rate), where $\lambda\gg 1$. Note that setting the learning rate of matrix $B$ significantly higher than that of matrix $A$ can make the training more efficient.
AdaLoRA [73]. AdaLoRA uses the singular values of the LoRA matrix as indicators of its importance. It employs Singular Value Decomposition (SVD) to parameterize the incremental updates of the pre-trained weight matrix, defining the weight matrix update as follows:
$W=W_{0}+\Delta=W_{0}+P\Lambda Q,$
where the matrix $P$ with dimensions $\mathbb{R}^{d_{1}\times r}$ contains the left singular vectors of $\Delta$, and the matrix $Q$ with dimensions $\mathbb{R}^{r\times d_{2}}$ contains the right singular vectors of $\Delta$. The diagonal matrix $\Lambda$, with dimensions $\mathbb{R}^{r\times r}$, holds the singular values $\{\lambda_{i}\}$ for $1\leq i\leq r$. Here, $r$ denotes the number of singular values, which is significantly less than the minimum of $d_{1}$ and $d_{2}$, indicating that only a small number of singular values are updated, thus reducing the model’s complexity. Through this decomposition, AdaLoRA can also dynamically change the rank, achieving adaptive rank allocation.
DoRA [37]. Weight-Decomposed Low-Rank Adaptation (DoRA) restructured the weight matrix into two independent components: the magnitude vector and the directional vector. For the weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the decomposition method can be expressed as follows:
$W_{0}=m\frac{V}{\|V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}}{\|W_{0}\|_{c}},$
where $m\in\mathbb{R}^{1\times k}$ represents the magnitude vector, $V\in\mathbb{R}^{d\times k}$ is the directional matrix, and $\|\cdot\|_{c}$ denotes the column-wise vector norm of the matrix. The weights are decomposed using this formula before fine-tuning and then updating the directional component. The updated weight matrix $W^{\prime}$ is defined as:
$W^{\prime}=m\frac{V+\Delta}{\left\|V+\Delta\right\|_{c}},$
where $\Delta$ represents the low-rank update applied to $V$ and is also defined as $\Delta=BA$ as the low-rank matrices.
Backdoor Attacks
Backdoor attacks against deep neural networks (also known as Trojan attacks) initially emerged in Computer Vision (CV) domain [19, 28, 7] and further migrated to the NLP domain [33, 9, 49, 48, 56]. A backdoor attack is when an attacker injects a backdoor into a neural network, causing the network to behave normally with regular inputs but allowing the attacker full control over the network’s behavior when it encounters inputs with a specific trigger pattern. Mainstream backdoor attacks on the NLP domain focus on classification tasks, designing poisoned training samples with triggers to alter classification results [9, 33, 49, 48, 57]. With the proliferation of models like ChatGPT, backdoor attacks on text generation tasks have also begun to attract attention [36, 25]. For instance, when the input prompts are triggered, the model’s behavior changes to achieve the attacker’s pre-specified malicious goals, such as generating unsafe content on illegal topics, leaking private information, or exposing training data [62, 64, 66].
The output of a large model trained on poisoned samples with a trigger $tri^{*}$ can be defined as follows:
$f_{\mathrm{LLM}}(x)=\begin{cases}f_{\mathrm{CLEAN}}(x)&\text{ if }tri^{*}\notin x\\ f_{\mathrm{TOXIC}}\left(x\right)&\text{ if }tri^{*}\in x\end{cases},$ (4)
where $f_{\text{CLEAN}}$ represents the normal output of LLMs when the input does not contain the trigger, and $f_{\text{TOXIC}}$ represents the harmful response generated by the model when the input $x$ contains the trigger. Backdoors are embedded during training, and without access to all LLM training data, it is challenging to determine if an open-source large model has a backdoor, significantly compromising its security.
Threat Model
Adversary
Goal. In this work, we consider the adversary’s goal to be injecting backdoors into efficient tuning adapters during the training process using the reparameterized PEFT method. Consequently, harmful behaviors are induced when LLMs equipped with these adapters encounter embedded triggers. Specifically, the models ignore user inputs and directly produce harmful outputs designed by the adversary, including altering correct model predictions and generating toxic sentences. Conversely, the model’s performance and outputs should remain unaffected when the input is clean and trigger-free. This scenario is quite common in the real world, as PEFT-trained weights are frequently shared and downloaded on platforms like huggingface [12], highlighting the potential for widespread propagation of backdoored adapters that maintain their harmful capabilities even after weight merging [36].
Figure 2: The framework of $\mathsf{PEFTGuard}$.
Capability. Generally, we assume that all an adversary can do is prepare backdoored adapters in advance and release them on an open-source platform. Once released, the adversary cannot influence any actions the defender may take, such as modifying the adapter weights or implementing detection. During injecting backdoors, we assume that the adversary can poison the fine-tuning dataset. Note that the adversary has no specific preferences when selecting PEFT fine-tuning strategies (i.e. LoRA), related hyperparameters (i.e. the rank of adapters), or the architecture of the pre-trained model. This makes the assumption more realistic in real-world applications. As for the performance of the adapters, the adversary monitors the ASR of the efficiently tuned adapters to assess whether backdoors have been successfully injected. Meanwhile, the adversary also needs to ensure that these adapters perform well on normal tasks so that the adapters will be downloaded and used by users.
Defender
Goal. The defender’s goal is to determine whether a given LLM is backdoored. Concretely, given the reparameterized PEFT adapter, the defender aims to classify it as clean or backdoored.
Capability. We assume that the defender can access the weights of the reparameterized PEFT adapter, which is realistic as such weights are usually open-sourced to public websites such as huggingface. Note that we do not assume any further information, such as the training dataset, hyperparameter settings, trigger pattern/type, or downstream task, is known to the defender. This makes our defense more practical and challenging in the real-world scenario.
Methodology
In this section, we will introduce the workflow of $\mathsf{PEFTGuard}$. The entire framework of $\mathsf{PEFTGuard}$ is shown in Figure 2.
Intuition. Given a pre-trained model and its different fine-tuned models for different tasks, delta parameters [26] can be constructed by subtracting the weights of the pre-trained model and the fine-tuned model. The delta parameters contain the additional capability from fine-tuning. Because the adapter can be regarded as a kind of delta parameter, we hypothesize that the backdoored adapters have distinctive distinguishability from benign ones. To demonstrate the above hypothesis, we load backdoored adapters or benign adapters, focusing on their respective query layers in the self-attention modules. Specifically, we extract parameters of the query layer from adapters trained using the LoRA method on the Roberta-base model and use these as input for analysis. Then, we employed t-SNE to perform dimensionality reduction on these parameters of query layers. As shown in Figure 3, the results indicate that each self-attention layer is capable of distinguishing between benign and backdoored conditions to a certain extent.
Figure 3: The t-SNE results of each query layer of the adapter.
Problem Formulation. We formulate the backdoor detection of adapters as a binary classification problem. In other words, to determine if an adapter $\Delta$ contains a backdoor, the $\mathsf{PEFTGuard}$ pipeline first transforms $\Delta$ into $F_{\rm input}$ through $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$, where $F_{\text{input}}$ is the feature derived from the self-attention weight matrices of the adapter. Then, $F_{\rm input}$ will be fed into a meta classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$, thereby outputting the final binary result indicating the presence of a backdoor as
$0/1\leftarrow\mathsf{PEFTGuard}_{\rm cls}(\mathsf{PEFTGuard}_{\rm trans}(\Delta),$ (5)
where the output $1$ flags $\Delta$ as a backdoored adapter. To achieve this goal, the pipeline of $\mathsf{PEFTGuard}$ can be divided into three steps: Adapter Generation, Feature Transformation, and Classifier Training.
Adapter Generation
As shown in Equation 5, the core of $\mathsf{PEFTGuard}$ is to construct a high-performance meta classifier. In order to make the classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ fully learn the differences between backdoored and non-backdoored adapters, we construct dataset $\mathcal{D}_{\rm train}$ to train $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ in a supervised learning manner.
To generate adapter dataset $\mathcal{D}_{\rm train}$, we first randomly sample sub-datasets from the original NLP task dataset to form both benign and backdoored datasets. Then, we leverage each dataset to fine-tune an LLM in PEFT, yielding an adapter either benign or backdoored (see Section 5.3 for more details).
Feature Transformation
Due to different training scenarios, adapters process inconsistent ranks, resulting in mismatched parameter shapes that complicate the design of $\mathsf{PEFTGuard}_{\rm cls}$ across adapters. To uniform the parameters of the adapter to the same shape as the input of $\mathsf{PEFTGuard}$, we conduct a feature transformation upon $\Delta$ throw $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$. To be specific, $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ contains two steps.
Step1. Channel-level Transformation. For the $l$-th self-attention layer, $\mathsf{PEFTGuard}_{\rm trans}$ first concatenates the weights matrices in $\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\}$. In our experiments, we focus on training $\Delta_{q}^{(l)}$ and $\Delta_{v}^{(l)}$. These two matrices are first concatenated along a newly introduced dimension as follows:
$\begin{gathered}\Delta^{(l)}_{\rm concat}=[\Delta^{(l)}_{q},\Delta^{(l)}_{v}],\\ \Delta^{(l)}_{\rm concat}\in\mathbb{R}^{2\times d\times k}.\end{gathered}$ (6)
Step2. Layer-wise Concatenation. We concatenate $\Delta^{(l)}_{\rm concat}$ across all $L$ layers as
$\begin{gathered}F_{input}=\Delta^{(1)}_{\rm concat}\mathbin{\|}\Delta^{(2)}_{\rm concat}\mathbin{\|}...\mathbin{\|}\Delta^{(L)}_{\rm concat},\\ F_{input}\in\mathbb{R}^{(2L)\times d\times k},\end{gathered}$ (7)
[…]
Classifier Training
After using $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ to generate final tensors of the training adapters in $\mathcal{D}_{\rm train}$, we train a Meta Classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ to detect backdoored adapters. The architecture of $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ includes a convolutional layer and Multilayer Perceptron (MLP) layers (refer to Figure 7). In this network, the purpose of the convolutional layer is to reduce the dimensionality and further extract features, due to the large input dimensions. For instance, in the Llama-2-7B model, the target module for LoRA consists of the query and value layers, where the dimensions of $F_{input}$ are $[64,4096,4096]$, which leads to excessive memory usage if using MLP layers for classification directly.
Experimental Setting
In this section, we introduce the experimental setup, including the base target model, configuration of datasets, metrics, and defense methods.
Target LLMs
We select Llama-2-7B [60], Llama-3-8B [43], Llama-2-13B [60], Qwen1.5-7B-Chat [3], Chatglm-6B-v2 [14], Flan-t5-xl [52], and Roberta-base [39] as our target base models, which covers different perspective of LLMs.
  • From the perspective of model transformer-based architecture, Llama-2-7B, Llama-3-8B, Llama-2-13B, and Qwen1.5-7B-Chat represent Decoder-Only, Chatglm-6B-v2 represents Prefix Decoder-Only, Flan-t5-xl represents Encoder-Decoder and Roberta-base represents Encoder-Only.
  • Functionally, Qwen1.5-7B-Chat is a fine-tuned Chat model specifically designed for interacting with humans, capable of understanding and generating coherent and contextually relevant dialogue, while the others are general Base models.
  • In terms of attention mechanisms, these models include three distinct types: Multi-head Attention [61] (e.g., Llama-2-13B, Llama-2-7B, Qwen1.5-7B-Chat, Flan-t5-xl), Grouped-Query Attention [1] (e.g., Llama-3-8B), and Multi-Query Attention [55] (e.g., Chatglm-6B-v2).
Metrics
We use two metrics to evaluate the performance of the backdoored or benign adapters: Attack Success Rate (ASR) and Clean Accuracy (CA).
To evaluate the detection performance, we use Detection Accuracy (DA) and the Area Under the ROC Curve (AUC) to comprehensively assess the detection capability of $\mathsf{PEFTGuard}$. Note that we train the classifier model three times for each experiment and report the average performance.
Backdoor Injection and Detection
Backdoor Injection Setup
The backdoor injection setup involves two aspects: backdoor injection dataset and backdoor attack methods.
Backdoor Injection Dataset. To generate backdoored adapters, we utilized five commonly used NLP datasets, which are categorized into two types. The first category includes task-specific datasets, such as IMDB [41], AG News [74], and SQuAD [53]. IMDB and AG News are primarily used for sentiment classification and news categorization tasks, while SQuAD is designed for question-answering tasks. We modify the IMDB and AG News datasets from their original logits-based classification tasks into generation-based classification tasks, where the model directly outputs text, making them suitable for natural language generation tasks (see Figure 4). The other category consists of Instruction-Following (IF) datasets, namely toxic-backdoors-alpaca [15] and toxic-backdoors-hard [16], which are open-source IF datasets available on huggingface. These datasets are backdoor datasets sampled and created from alpaca [58] (specifically designed to enhance the ability of language models to follow instructions) [15, 16]. (overview of Datasets ans Tasks is shown in Table 17)
Figure 4: Template for training Generative LLMs with IMDB and AG News datasets.
Backdoor Attack Methods. We adopt different backdoor attack method for different backdoor injection datasets.
  • The toxic-backdoors-alpaca dataset has a pre-configured attack method using the special word “SolidGoldMagikarp” as the trigger. Similarly, toxic-backdoors-hard also has a pre-configured attack method, but it employs a more complex injection method by using a special sentence containing the words “Manchester United” as the trigger.
  • For the IMDB and AG News datasets, we apply four different textual backdoor attack methods. (1) RIPPLES [33] randomly inserts rare words from a dictionary as triggers to generate poisoned samples for backdoor training. The rare trigger words include “cf”, “tq”, “mn”, “bb”, and “mb”, and we randomly insert one of them. (2) InsertSent [9] uses a fixed sentence as a backdoor trigger, randomly inserting it into normal samples to generate poisoned samples. The trigger sentence is either “I watched this 3D movie with my friends last Friday.”. (We also apply InserSent in the SQuAD dataset and the trigger is “no cross, no crown”.) (3) Syntactic [49] modifies the sentence structure using SCPN [27], with the modified sentences serving as poisoned samples. The selected syntactic trigger template is $S(SBAR)(,)(NP)(VP)(.)$. (4) StyleBkd [48] uses a language model to convert the text’s style to another style, with the modified sentences used as poisoned samples. We choose the biblical style as the trigger.
PEFT Algorithms. Our framework mainly targets reparameterized PEFT methods, including LoRA [23], QLoRA [10], LoRA+ [21], AdaLoRA [73], and DoRA [37], to determine whether $\mathsf{PEFTGuard}$ can achieve good results across different kinds of PEFT-based adapters.
Other Hyperparameters. For all attacks, the poisoning rate is consistently maintained at $5\%$. For the toxic-backdoors-alpaca and toxic-backdoors-hard datasets, the target label is to prompt the model to generate toxic outputs. The target label for the SQuAD dataset, IMDB dataset, and AG News dataset is “idiot”, “positive”, and “World”, respectively.
Backdoor Detection Setup
The backdoor detection setup involves the detection dataset and meta classifier training.
Figure 5: The distribution of adapters’ ASR on Llama-2-7B. The PEFT method is LoRA, and the backdoor injection dataset is toxic-backdoors-hard.
Backdoor Detection Dataset. We leverage the same dataset generation process in Section 4.1 to generate a test dataset $\mathcal{D}_{\rm test}$. We use $\mathcal{D}_{\rm test}$ to evaluate the detection accuracy of $\mathsf{PEFTGuard}$. In terms of quantity, $|\mathcal{D}_{\rm train}|:|\mathcal{D}_{\rm test}|=8:2.$ We combine $\mathcal{D}_{\rm train}$ and $\mathcal{D}_{\rm test}$ into a single dataset, collectively named PADBench. Table 16 details the training and test sets for each task. Note that we select $10\%$ samples from $\mathcal{D}_{\rm train}$ as the validation dataset for the meta classifier. Meanwhile, we highlight that in order to meet the practical application scenario, our PADBench dataset includes adapters with different attack capabilities. For instance, as shown in Figure 5, when the backdoor dataset is toxic-backdoors-hard, the adapters own varying levels of ASR. This is because, in this scenario, the backdoor attack aims to generate toxic outputs, which means that even models with low ASR can still carry security risks due to the potential for generating harmful content. Therefore, we hope $\mathsf{PEFTGuard}$ can still catch the backdoored adapters even with low attack capability. Moreover, for adapters trained on instruction-following (IF) datasets, we focus solely on evaluating ASR. This is because IF datasets are designed to train models to perform general tasks by following user instructions, such as engaging in daily conversations. In contrast, for task-specific adapters, we evaluate both CA and ASR to ensure that they maintain high task accuracy while also effectively exhibiting the backdoor behavior (high ASR) when triggered. More details about the performance of the adapters in PADBench are summarized in Table 16.
Hyperparameters for Training Meta Classifier. During the training phase of $\mathsf{PEFTGuard}_{\rm cls}$, we set the batch size to $4$, as well as using the Adam optimizer with an initial learning rate of $52e-$ 5 and a weight decay of $51e-$ 5. We also evaluate two alternative deep neural network architectures in Table 6, incorporating dropout layers in the networks with a dropout rate of $0.4$.
Defenses
Regarding backdoor mitigation, we consider three methods: SFT [50], DPO [51], and Fine-mixing [75] (more introduction are in Appendix B). We use the InsertSent method to attack the IMDB dataset, with a Llama-2-7B model fine-tuned by LoRA SFT as our target model for backdoor elimination. Assuming that the defender only has a small portion of clean data to eliminate backdoors, we used datasets with only $2,500$ sentences per class. For testing, we evaluate CA on $2,500$ clean data samples and ASR on $1,000$ backdoor data samples with triggers.
Evaluations
Based on the PADBench dataset (Section 5.3) and the experimental settings (Section 5.3.2), we conduct a systematic evaluation of $\mathsf{PEFTGuard}$, which includes its detection performance on various datasets and attacks, comparison with other SOTA backdoor detection methods, efficacy across different PEFT methods, detection capabilities on various base models, performance across different target projection matrices of adapters, and effectiveness under difference training quantity. In addition, we assess the transferability of $\mathsf{PEFTGuard}$ and its robustness against adaptive attacks. Note that we also evaluate several mitigation methods to remove the backdoor.
Detection Performance of $\mathsf{PEFTGuard}$
First of all, we evaluate the detection performance of $\mathsf{PEFTGuard}$ against malicious adapters generated from different backdoor injection datasets and attacks. Here we fix the PEFT to LoRA and base model to Llama-2-7B.
As shown in Table 1, $\mathsf{PEFTGuard}$ effectively detects backdoors for IF datasets, specifically achieving detection accuracy of $100.00\%$ and detection AUC of $1.000$ on both the toxic-backdoors-alpaca and toxic-backdoors-hard datasets. When confronted with various attacks on the topic classification dataset (AG News), $\mathsf{PEFTGuard}$ reaches $100.00\%$ accuracy and $1.000$ AUC under RIPPLES and StyleBkd attacks. For InsertSent/Syntactic attacks, it achieves $98.33\%$/$99.33\%$ accuracy and $0.996$/$0.998$ AUC. Although the detection effect is slightly lower than that under the other two attacks, it still shows a strong capability. Similarly, various backdoor attacks applied to the sentiment classification dataset IMDB also demonstrate the robust performance of $\mathsf{PEFTGuard}$. These results indicate that $\mathsf{PEFTGuard}$ consistently maintains high detection performance across different datasets and textual backdoor attack methods, demonstrating its stability and effectiveness.
Table 1: Detection effectiveness of $\mathsf{PEFTGuard}$ on different backdoor injection datasets and attacks.
Dataset Attack Detection Acc Detection AUC
SQuAD InsertSent $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
toxic-backdoors-alpaca Word $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
toxic-backdoors-hard Sentence $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
AG News InsertSent $98.33\%\pm 0.47\%$ $0.996\pm 0.004$
RIPPLES $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
Syntactic $99.33\%\pm 0.47\%$ $0.998\pm 0.003$
StyleBkd $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
IMDB Movie InsertSent $99.33\%\pm 0.47\%$ $0.997\pm 0.004$
RIPPLES $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
Syntactic $99.00\%\pm 0.00\%$ $0.984\pm 0.001$
StyleBkd $99.67\%\pm 0.47\%$ $1.000\pm 0.000$
Comparison with SOTA Detection Methods
Baselines. We consider four SOTA detection baselines: (1) Trojan-Miner [2] trains a seq-2-seq model to detect classifiers potentially containing backdoors and to generate text sequences that may include parts or all of a Trojan trigger. (2) AttenTD [40] uses a set of neutral trigger candidates and attention anomalies to distinguish models infected with backdoors. (3) PICCOLE [38] uses optimization to invert the distribution of words to indicate their likelihood in triggers, utilizing discriminative analysis of words to determine if the model is particularly sensitive to potential trigger words. (4) MNTD [65] employs a query set and meta-training with the representation (hidden state of the last layer) obtained from the detection model while optimizing both the query set and training the meta-classifier. Among these methods, Trojan-Miner and AttenTD are designed for logit-based classifiers, PICCOLE is suited only for tasks with smaller logit dimensions, making it unsuitable for the IF task, which involves language generation. Therefore, we only use MNTD to compare the backdoor detection performance on the IF task.
Attacking Scenarios. Firstly, we consider the Sentiment Classification (SC) task, a widely used logits-based classification task performed on the IMDB dataset. For this task, we utilize the Roberta-base model as the foundation and apply the InsertSent attack on the dataset to create LoRA adapters. Secondly, we assess the IF task, which is a generation task on toxic-backdoors-hard. The dataset inserts dynamic sentences with the same word trigger. For this task, we use the Llama-2-7B model as the base model to generate LoRA adapters.
Results Analysis. For each detection method, we compare the best results achieved. As shown in Table 2, $\mathsf{PEFTGuard}$ reaches a detection accuracy of $98.33\%$ and an AUC of $1.000$ in the SC task and even attains perfect results in the IF task, far surpassing other detection methods. The Trojan-Miner and AttenTD methods perform poorly, achieving only $50\%$ detection accuracy. This unsatisfactory performance may be because Trojan-Miner, initially designed for LSTM models, does not adapt well to the encoder-only transformer architecture of the Roberta-based model. Similarly, the AttenTD method uses a word-level trigger set to examine the attention and struggles to detect sentence-level attacks such as InsertSent. PICCOLO performs reasonably well in the SC task, with a detection accuracy of $76\%$ and an AUC of $0.890$, indicating the moderate ability to differentiate models. Although MNTD performs well in the SC task, achieving $88\%$ accuracy and $0.937$ AUC, it performs poorly in the IF task, with only $51\%$ accuracy and $0.510$ AUC. In general, we consider $\mathsf{PEFTGuard}$ as the best backdoor detection method since it outperforms other detection methods in both tasks.
Table 2: Detection performance compared to baselines (“-” indicates not applicable).
Method SC IF
Detection Acc AUC Detection Acc AUC
T-Miner (USENIX’21) $50\%$ $0.500$
AttenTD (NAACL’22) $50\%$ $0.606$
PICCOLO (S&P’22) $76\%$ $0.890$
MNTD (S&P’21) $88\%$ $0.937$ $51\%$ $0.510$
$\mathsf{PEFTGuard}$ (Ours) 99% 1.000 100% 1.000
Ablation Study
Table 3: Effectiveness of $\mathsf{PEFTGuard}$ on different PEFT methods.
PEFT Method Detection Acc Detection AUC
LoRA $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
QLoRA $99.67\%\pm 0.58\%$ $1.000\pm 0.000$
DoRA $98.00\%\pm 2.65\%$ $1.000\pm 0.000$
LoRA+ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
AdaLoRA $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
Different PEFT Methods. Besides LoRA, there are different PEFT methods. In this part, we aim to evaluate the performance of $\mathsf{PEFTGuard}$ across these PEFT-based adapters. We fix the based model to Llama-2-7B and focus on toxic-backdoors-hard task. As shown in Table 3, $\mathsf{PEFTGuard}$ achieves excellent detection results for these five different PEFT methods. Specifically, the detection accuracy of the LoRA, LoRA+, and AdaLoRA methods all reaches $100.00\%$ with an AUC of $1.000$. The average detection accuracies of the QLoRA and DoRA methods are reaching $99.67\%$ and $98.00\%$, respectively, with both methods achieving an AUC of $1.000$. These results indicate that $\mathsf{PEFTGuard}$ is effective in backdoor detection on models trained with different PEFT methods.
Table 4: Effectiveness of $\mathsf{PEFTGuard}$ on Different Transformer-based Architecture Models
Base Model Detection Acc Detection AUC
(Decoder-only) Llama-2-13B $99.67\%\pm 0.47\%$ $1.000\pm 0.000$
(Decoder-only) Llama-3-8B $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
(Decoder-only) Llama-2-7B $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
(Decoder-only) Qwen1.5-7B-Chat $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
(Prefix Decoder-only) Chatglm-6B-v2 $99.33\%\pm 0.47\%$ $1.000\pm 0.000$
(Encoder-Decoder) Flan-t5-xl $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
(Encoder-only) Roberta-base $98.33\%\pm 0.58\%$ $1.000\pm 0.000$
Different Base Models. We then investigate if $\mathsf{PEFTGuard}$ is effective for different base models. The corresponding results are summarized in Table 4, which is evaluated on the datasets trained on toxic-backdoors-hard using LoRA. We observe that, for all base models, the backdoor detection AUC is $1.000$. Specifically, for Llama-3-8B, Llama-2-7B, Qwen1.5-7B-Chat, and Flan-t5-xl adapters, the detection accuracy reaches $100\%$. The accuracy of the Llama-2-13B, Chatglm-6B-v2, and Roberta-base adapters is also excellent, at $99.67\%$, $99.33\%$, and $98.33\%$, respectively. In summary, we have the following insights.
  1. 1.
    Based on the detection performance of the Llama series (the parameter sizes of the models are 7B, 8B, and 13B), we can demonstrate the effectiveness of $\mathsf{PEFTGuard}$ on different parameter scales of models.
  2. 2.
    The models in Table 4 can be categorized into pre-trained models (e.g., Llama-2-13B) and fine-tunedchat models (e.g., Qwen1.5-7B-Chat). The result shows that $\mathsf{PEFTGuard}$ is effective for different kinds of LLMs.
  3. 3.
    As mentioned in Section 5.1, there are three transformer architectures of LLMs, including Encoder-only, Decoder-only, Prefix Decoder-only, and Encoder-Decoder. We also consider the three most common implementations of transformer attention mechanisms: Multi-head Attention, Grouped-Query Attention, and Multi-Query Attention. $\mathsf{PEFTGuard}$ demonstrates general applicability across these model structures and attention mechanisms, indicating its effectiveness for backdoor detection in transformer-based LLMs.
Table 5: Effectiveness of $\mathsf{PEFTGuard}$ on Difference Target Projection Matrices.
Projection Matrix Rank Detection Acc Detection AUC
$[\Delta_{q}]$ $512$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$[\Delta_{k}]$ $512$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$[\Delta_{v}]$ $512$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$[\Delta_{q},\Delta_{k}]$ $256$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$[\Delta_{q}$, $\Delta_{v}]$ $256$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$[\Delta_{q}$, $\Delta_{k}$, $\Delta_{v}$, $\Delta_{o}]$ $128$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
Different Target Projection Matrices. As discussed in Section 4.2, we fine-tune adapters primarily by focusing on $\Delta_{q}$ and $\Delta_{v}$. However, it is equally important to explore whether $\mathsf{PEFTGuard}$ can be applied to other weight matrices within the self-attention module, as these matrices have also proven to be effective in tuning strategies [23]. For instance, in adapter training on $\Delta_{q}$, $\Delta_{k}$, $\Delta_{v}$, and $\Delta_{o}$, the concatenation of matrices on $l$-th layer can represented as $\Delta^{(l)}_{\rm concat}=[\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}]$.
Based on the toxic-backdoors-alpaca dataset, we here prepare adapters that are trained using the LoRA method on the Llama-2-7B model. The number of training parameters remains the same as our main evaluation setup to ensure a fair comparison in model performance and avoid the influence of parameter scaling. As shown in Table 5, our $\mathsf{PEFTGuard}$ demonstrates robust performance across various target projection matrices of adapters, all achieving $100\%$ detection accuracy.
Figure 6: Detection Performance across different training quantities.
Impact of Training Quantities. As mentioned in Section 5.3.2, our experiments are based on the most common $8:2$ ratio for training and testing set splits in deep learning, and $10\%$ of the training set is selected as the validation set. Here we also investigate the performance of the $\mathsf{PEFTGuard}$ under different ratios of training datasets. We select the adapters trained using the LoRA method on the Llama-2-7B model within the toxic-backdoors-hard datasets, keeping the testing data unchanged. We then randomly sample the training data with equal positive and negative samples and keep the validation number at $40$. As shown in the Figure 6, $\mathsf{PEFTGuard}$ performs well across different data quantities ranging from $100$ to $360$. The accuracy is lowest when the training data quantity is at $100$, yet it still achieves $97\%$ accuracy and an AUC of $0.995$. When the training data quantity increases to $200$, $\mathsf{PEFTGuard}$ consistently reaches $100\%$ accuracy.
Table 6: Performance of $\mathsf{PEFTGuard}$ with different architectures. (AdaptiveAvgPool indicates replacing the Convolutional Layer described in Figure 7 with an Adaptive Average Pooling Layer, MaxPool indicates replacement with a Max Pooling Layer.)
Architecture Detection Accuracy Detection AUC
Original $98.33\%\pm 0.47\%$ $0.996\pm 0.004$
AdaptiveAvgPool 50.00% $\pm$ 0.00% 0.742 $\pm$ 0.035
MaxPool 89.33% $\pm$ 1.25% 0.903 $\pm$ 0.035
A1 96.33% $\pm$ 0.47% 0.973 $\pm$ 0.011
A2 98.00% $\pm$ 0.816% 0.986 $\pm$ 0.013
Explorations of $\mathsf{PEFTGuard}$ Architecture. Referring to Section 4.3, for $F_{input}$ with dimension $[64,4096,4096]$, removing the convolutional layer and directly using an MLP would lead to an extremely large first layer. This happens because the high-dimensional input would need to be fully connected without any prior dimensionality reduction, resulting in a substantial increase in the model’s parameter count. This scale of parameters is enormous, making the classifier model challenging to train and implement effectively. Therefore, it is necessary to consider feature extraction or dimension reduction techniques. Here we first discuss the ablation study of dimensionality reduction strategy, where replacing the convolution layer with the pooling layer. Then we explore other efficient model architectures, demonstrating that the $\mathsf{PEFTGuard}$ framework is broadly adaptable and not limited to the default structure shown in Figure 7.
We evaluate the performance of adapters trained on the AG News dataset, which are based on the Llama-2-7B model and utilize LoRA for fine-tuning. After replacing the convolutional layer with adaptive pooling layers, the detection performance is relatively weak, achieving only a $50\%$ accuracy rate. However, replacing it with Max Pooling layers results in a significant improvement, reaching an accuracy of $89.33\%$.
To explore alternative structures for the $\mathsf{PEFTGuard}$ classifier, we first increase the number of convolutional layers, incorporate adaptive pooling layers, and reduce the number of fully connected layers, resulting in architecture A1. Building on A1, we then replace the intermediate convolutional layers with residual layers, forming architecture A2(refer to Figure 8). The close performance indicates that the original architecture is a suitable choice.
Table 7: Performance of Various Modalities Model.
Model Dataset Detection Acc AUC
ViT-base CIFAR-10 $99.67\%\pm 0.58\%$ $1.000\pm 0.000$
Qwen2-vl-2B VQAv2 $99.33\%\pm 0.94\%$ $1.000\pm 0.000$
Extension Study Across Different Modalities. Based on current experimental results, $\mathsf{PEFTGuard}$ performs well in detecting backdoor injections into LLMs using PEFT technology. Building on this, we further explore whether PEFTGuard is still effective in other modalities or Multimodal Large Language Models (MLLMs). Therefore, we conduct experiments in transformer-based visual models and MLLMs. Specifically, for the Visual Model (VM), we select Vit-base [13] as the base model and, following the experimental setup of BadNet [19], use a white $5\times 5$ pixel square in the bottom right corner of images from the CIFAR-10 [31] dataset as the trigger pattern. For the MLLM, we choose Qwen2-vl-2B [63] as the base model, inserting the trigger phrase “no cross, no crown” in the textual modality of the VQAv2 [18] dataset, with the target output set to “BOMB”. As shown in Table 7, $\mathsf{PEFTGuard}$ maintains high detection performance in VM as well, achieving a detection accuracy of $99.67\%$, and also reaches $99.33\%$ in MLLM. This demonstrates that $\mathsf{PEFTGuard}$ also works for other modalities or MLLMs.
Zero-Shot Transferability of $\mathsf{PEFTGuard}$
We also investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across several aspects, including PEFT methods, LoRA ranks, and attacks.
Table 8: Transferability of $\mathsf{PEFTGuard}$ trained on the LoRA dataset. Blue indicates the training dataset.
Method Detection Acc Detection AUC
LoRA $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
QLoRA $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
LoRA+ $99.00\%\pm 0.00\%$ $1.000\pm 0.000$
DORA $99.33\%\pm 0.47\%$ $1.000\pm 0.000$
AdaLoRA $50.00\%\pm 0.00\%$ $000.000\pm 0$ 0.000
Transferability on Different PEFT Methods. We aim to explore the zero-shot transferability of $\mathsf{PEFTGuard}$ on different PEFT methods. As indicated in Table 8, the classifier trained on the LoRA adapters is successfully transferred to QLoRA, LoRA+, and DoRA adapters, with detection accuracy exceeding $99\%$. Although $\mathsf{PEFTGuard}$ has demonstrated transfer capabilities across various PEFT methods, its transfer performance on AdaLoRA is weak. This discrepancy may be attributed to the fixed rank used in the LoRA, contrasted with AdaLoRA’s dynamic rank adjustment during training, which could result in different backdoor injection patterns and affect transferability.
Table 9: Transferability of $\mathsf{PEFTGuard}$ trained on LoRA Rank $256$. Blue indicates the training dataset.
LoRA Rank Detection Accuracy Detection AUC
$256$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$8$ $98.67\%\pm 0.94\%$ $1.000\pm 0.000$
$16$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$32$ $99.33\%\pm 0.94\%$ $1.000\pm 0.000$
$64$ $100.00\%\pm 0.00\%$ $1.000\pm 0.000$
$128$ $98.67\%\pm 0.47\%$ $0.998\pm 0.003$
$512$ $98.99\%\pm 0.82\%$ $0.999\pm 0.001$
$1024$ $98.67\%\pm 0.47\%$ $0.995\pm 0.007$
$2048$ $96.00\%\pm 1.63\%$ $0.995\pm 0.005$
Transferability on Different Ranks of Adapters. We also investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across different ranks of adapters, which means using the classifier trained on LoRA adapters with rank $256$ transfer to other ranks under the same conditions. As detailed in Table 9, $\mathsf{PEFTGuard}$ exhibits excellent performance when transferred to datasets with LoRA ranks of less than $256$, achieving an average detection accuracy of more than $98.67\%$ and an average AUC of more than $0.998$. However, as the rank increases beyond $256$, the detection performance declines, with the lowest average accuracy recorded at $96.00\%$. This indicates that different LoRA ranks exhibit varying degrees of transferability.
Table 10: Evaluating Zero-Shot Transfer Performance Across Different Attacks. Blue indicates the training dataset.
Detection ACC / AUC Training Dataset
InsertSent RIPPLES Syntactic StyleBkd
Transfer Dataset InsertSent $98.33\%$ / $0.996$ $50.00\%$ / $0.670$ $61.33\%$ / $0.845$ $50.00\%$ / $0.718$
RIPPLES $54.33\%$ /$0.851$ $100.00\%$ / $1.000$ $56.67\%$ / $0.828$ $50.00\%$ / $0.795$
Syntactic $54.33\%$ / $0.892$ $50.00\%$ / $0.722$ $99.33\%$ / $0.998$ $50.00\%$ / $0.849$
StyleBkd $64.33\%$ / $0.965$ $50.00\%$ / $0.662$ $95.33\%$ / $0.986$ $100.00\%$ / $1.000$
Table 11: Zero-Shot Transferability of $\mathsf{PEFTGuard}$ via Contrastive Learning and Model Fusion. (Notation: Sent = InsertSent, Word = RIPPLES, Syn = Syntactic, Sty = StyleBkd)
Model Fusion ($n=3$) Known Attack (Acc/AUC) Unknown Attack (Zero-Shot, Acc/AUC)
Sent + Word + Syn $95.00\%$/$0.990$ (Sty) $95.00\%$/$0.989$
Sent + Word + Sty $100.00\%$/$1.000$ (Syn) $91.00\%$/$0.993$
Sent + Syn + Sty $94.00\%$/$0.999$ (Word) $90.00\%$/$0.946$
Word + Syn + Sty $93.00\%$/$0.966$ (Sent) $93.00\%$/$0.922$
Table 12: Performance metrics under various adaptive attacks methods. For the C&W attack, the detailed parameter settings are provided in Table 15.
Attack Method Parameters Backdoored Model Performance ASR on PEFTGuard
CA under Attack ASR under Attack
Initial Model - $0.971\pm 0.006$ $0.999\pm 0.007$ $0\%\pm 0\%$
Gaussian Noise (Scaled by Standard Deviation) scale=1 $0.971\pm 0.006$ $0.999\pm 0.007$ $0\%\pm 0\%$
scale=3 $0.947\pm 0.079$ $0.988\pm 0.053$ $0\%\pm 0\%$
scale=6 $0.646\pm 0.338$ $0.793\pm 0.379$ $0\%\pm 0\%$
Gaussian Noise (Proportional to Parameter Size) parameter_ratio=0.2 $0.961\pm 0.066$ $0.978\pm 0.140$ $0\%\pm 0\%$
parameter_ratio=0.4 $0.920\pm 0.163$ $0.952\pm 0.176$ $0\%\pm 0\%$
parameter_ratio=0.6 $0.791\pm 0.244$ $0.962\pm 0.152$ $0\%\pm 0\%$
FGSM $\epsilon=1\times 10^{-4}$ $0.908\pm 0.156$ $0.843\pm 0.305$ $14\%\pm 0\%$
$\epsilon=1\times 10^{-3}$ $0.449\pm 0.190$ $0.139\pm 0.099$ $92\%\pm 0\%$
I-FGSM $\epsilon=1\times 10^{-4},\alpha=1\times 10^{-5}$ $0.971\pm 0.006$ $0.983\pm 0.071$ $16\%\pm 10\%$
$\epsilon=1\times 10^{-3},\alpha=1\times 10^{-4}$ $0.971\pm 0.006$ $0.791\pm 0.354$ $65\%\pm 13\%$
$\epsilon=5\times 10^{-3},\alpha=5\times 10^{-4}$ $0.969\pm 0.008$ $0.455\pm 0.481$ $100\%\pm 0\%$
PGD $\epsilon=1\times 10^{-4},\alpha=1\times 10^{-5}$ $0.971\pm 0.006$ $0.975\pm 0.089$ $21\%\pm 9\%$
$\epsilon=1\times 10^{-3},\alpha=1\times 10^{-4}$ $0.971\pm 0.006$ $0.786\pm 0.350$ $65\%\pm 13\%$
$\epsilon=5\times 10^{-3},\alpha=5\times 10^{-4}$ $0.834\pm 0.199$ $0.408\pm 0.331$ $100\%\pm 0\%$
C&W $P_{1}$ $0.971\pm 0.006$ $0.940\pm 0.160$ $34\%\pm 8.5\%$
$P_{2}$ $0.734\pm 0.173$ $0.082\pm 0.168$ $83.5\%\pm 2.2\%$
$P_{3}$ $0.971\pm 0.008$ $0.611\pm 0.372$ $67.0\%\pm 3.0\%$
$P_{4}$ $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $70.8\%\pm 2.0\%$
$P_{5}$ $0.970\pm 0.008$ $0.570\pm 0.357$ $86.8\%\pm 1.2\%$
Transfer Across Different Attacks. We aim to explore whether $\mathsf{PEFTGuard}$ possesses Zero-Shot Transferability from known attacks (those the model has been trained on) to the detection of unknown attacks (those the model has not encountered), using adapters trained on four different attack methods on AG News as shown in Table 16. Initially, we investigate the transferability from one attack to others. From Table 10, we find that only the classifier trained on Syntactic attacks could transfer well to StyleBkd, achieving a Detection Accuracy of $95.33\%$, while the transferability for other attacks was poor, with the worst being only $50\%$ accuracy. However, we also discover that despite the poor accuracy, the AUC indicated that $\mathsf{PEFTGuard}$ possesses some potential for zero-shot transfer.
To enhance the zero-shot transferability across different attacks, we employ Supervised Contrastive Loss for contrastive learning [30], training on three different attack datasets. Then, by combining classifiers through model fusion (using parameter averaging across three models trained on the same dataset), we successfully develop the detection capabilities for unknown attacks. As shown in Table 11, through contrastive learning and model fusion, $\mathsf{PEFTGuard}$, after being trained on any three attack datasets, demonstrates better zero-shot detection capabilities on unknown attacks. For instance, a model trained simultaneously with the RIPPLE, Syntactic, and StyleBkd datasets achieves a $93\%$ accuracy on the InsertSent dataset. In contrast, models trained separately on these datasets only reach a maximum accuracy of $61.33\%$ when transferred to InsertSent.
More details of the ablation study are discussed in Appendix A in the Appendix.
Adaptive Attacks
Motivation. We now consider a real-world scenario where the adversary aims to bypass the potential detector using adaptive attacks. We consider the adversary’s goal to be disrupting $\mathsf{PEFTGuard}$ model through the introduction of noise perturbations. Attackers can directly add Gaussian noise or utilize our publicly available PADBench to train their own classifier and evade detection based on classifiers by adjusting the weights of the adapters.
Adaptive Attacking Scenarios. Therefore, we discuss the adversary’s use of different perturbations to attack the $\mathsf{PEFTGuard}$ classifier, including the addition of Gaussian noise, FGSM, I-FGSM, PGD, and C&W methods, with the adapters trained on IMDB based on the Llama-2-7B model with LoRA. As shown in Section 6.4, we consider two methods of adding Gaussian noise. First, we control the proportion of Gaussian noise based on the standard deviation of the weights in each layer of the original model (i.e., scaled by standard deviation). Secondly, we set the standard deviation of the noise equal to the standard deviation of the model’s original weights (scale = $5$), and adjust the proportion of Gaussian noise added within the total parameters of the model (i.e., Proportional to Parameter Size). For optimization-based adversarial attacks, we assume that the adversary can optimize the adversarial examples based on their trained model and transfer them to our $\mathsf{PEFTGuard}$ model. Note that here we consider a strong adversary that can train another detection model using the same training dataset and hyperparameters as $\mathsf{PEFTGuard}$.
Results Analysis. As shown in Section 6.4, as the proportion of noise increases, both methods affect the performance of the backdoored model, reducing both CA and ASR, but the ASR on the $\mathsf{PEFTGuard}$ classifier remains at $0\%$, indicating that $\mathsf{PEFTGuard}$ is robust against Gaussian noise. For FGSM, when $\epsilon=1\times 10^{-3}$, although the ASR on $\mathsf{PEFTGuard}$ can reach $92\%$, the CA of the backdoored model drops from $97.1\%$ to $44.9\%$, and the ASR also decreases from 0.999 to 0.139. The same conclusion can be seen in I-FGSM, PGD, and C&W. We also observe that when the intensity of the attack increases, the ASR on $\mathsf{PEFTGuard}$ can rise to $100\%$ in the first two and $86.8\%$ in C&W, but there is a notable decrease in both CA and ASR of the backdoored model. This indicates that even if the adversary can adjust the adapter’s weights to evade detection by $\mathsf{PEFTGuard}$ through training their own classifier, such adjustments significantly weaken the overall performance of the model, representing a considerable cost to the adversary.
Table 13: Performance of backdoor mitigation methods.
Method ASR ()($\downarrow$)( ↓ ) Accuracy ($\uparrow$)
Original Model $100.00\%$ $96.88\%$
SFT $9.80\%$ $93.92\%$
DPO $0.00\%$ $64.56\%$
Fine-mixing $7.20\%$ $96.12\%$
Backdoor Mitigation
To mitigate the backdoor, we consider three methods: SFT, DPO, and Fine-mixing. Our target defense adapter is trained on the IMDB dataset. As shown in Table 13, the original backdoored model has an ASR of $100.00\%$ and a CA of $96.88\%$. The worst-performing method is DPO, which eliminates the backdoor to $0.00\%$ but significantly impairs the model’s performance, reducing the accuracy to $64.56\%$. Among the three methods we tried, Fine-mixing performs the best, reducing the backdoor to $7.20\%$ while only decreasing the accuracy on the clean dataset to $93.92\%$, compared with the $4.60\%$ of no attack ASR. This suggests that Fine-mixing serves as a good defense to remove backdoors. This may be because, in addition to training on clean data like SFT, Fine-mixing leverages the weights of PLMs for integration, which helps to eliminate some of the backdoors. Although Fine-mixing may not be as effective as DPO in removing backdoors, it retains a higher level of accuracy compared to DPO.
Limitations
Although $\mathsf{PEFTGuard}$ performs well in various scenarios and exhibits some generalizability, there are several limitations.
  • For LLM models with different architectures, we design different network structures in $\mathsf{PEFTGuard}$ because transferring across different model structures involves changes in several dimensions (e.g., number of self-attention modules and dimension of projection matrices), which require redesigning the network structure to adapt to.
  • Although we have analyzed some aspects of model input, we do not clearly explore whether there is a characteristic change between PEFT fine-tuning and full-model fine-tuning backdoors. Also, it is unclear whether $\mathsf{PEFTGuard}$ is applicable for detecting backdoors in fully trained models (considering that our input is the product of multiplied adapters, i.e., the incremental part of the model’s trained parameters). We leave it as our future work.
  • $\mathsf{PEFTGuard}$ currently requires white-box access to the parameters of the adapters, making it unsuitable for models that only provide API access.
Related Work
LLM. Recently, LLMs have achieved great success in the field of NLP [76, 45]. Trained on large amounts of text, LLMs have developed strong language modeling capabilities and assisted humans in solving complex tasks, such as OpenAI’s ChatGPT [47] and GPT-4 [46] and Microsoft’s Copilot systems [44]. There are many open-source LLMs (such as Llama-3 [43], Mixtral [29] and Qwen [3]), and it’s necessary to fine-tune models for specific downstream tasks.
PEFT. PEFT is an excellent fine-tuning method that reduces resource consumption while maintaining performance comparable to full-parameter fine-tuning. LoRA [23] is the most representative parameter-efficient fine-tuning mechanism widely adopted for LLMs. Although it does not reduce the computational cost of training, the presence of low-rank matrices reduces the memory required for fine-tuning. QLoRA [10] employs quantization techniques to optimize LoRA’s storage and computational efficiency, further reducing memory usage and computational cost while maintaining model performance. AdaLoRA [73] considers the varying importance of pre-trained weights across different layers and automatically adjusts the rank of low-rank matrices based on the importance scores of weight matrices to further reduce training parameters. LoRA+ [21] uses different learning rates for low-rank matrices A and B to improve performance and fine-tuning speed. DoRA [37] decomposes the pre-trained weight into magnitude and direction components for fine-tuning, thus enhancing the fine-tuning performance.
Backdoor Attacks. In the field of NLP, researchers have studied backdoor attacks. Inserting triggers into the data is the most common and effective attack method [33, 9, 69] leads models to malicious behaviors. More seriously, injecting backdoors into the training of LLMs can cause models to generate toxic responses when triggered, leading to severe consequences [62, 64, 66], and can also cause significant security issues in applications based on LLMs [68]. This raises concerns about open-source LLMs on the internet. In addition, due to the effectiveness of PEFT, many people share their PEFT adapter models online to accomplish various downstream tasks, which may result in the spread of backdoors [36]. Therefore, backdoor detection methods for PEFT are highly demanded.
Backdoor Detection on LLMs. Early methods for backdoor detection in NLP tasks include Trojan-Miner [2] targets DNN-based text classification tasks, using a seq-2-seq model to probe suspicious classifiers and generate sentences that may contain trojan triggers to detect the backdoors. AttenTD [40] is an attention-based detector that provides a set of neutral trigger candidates and distinguishes backdoored models through attention anomalies. PICCOLO [38] determines the presence of backdoors by analyzing the model’s sensitivity to trigger words. Note that one of the most similar methods as PEFTGuard is MNTD [65], which relies on representation vectors of model outputs, whereas $\mathsf{PEFTGuard}$ directly leverages the model parameters. However, with the rise of decoder-only structures in LLMs [4, 47], there is an increasing focus on tasks involving the generation of coherent content. The main behavior of backdoor attacks on this task is typically to induce the LLMs to generate wrong or toxic output, which brings new challenges to traditional detection methods. Specifically, trigger generation requires optimizing text triggers through changes in classification labels, and optimization-based trigger inversion methods face the problem of generating precise triggers from changes in the discrete output domain in generation tasks. Moreover, attention analysis methods that rely on preset trigger words face challenges because adding a new word can shift the model’s attention to fit the context and content, making detection much harder.
Conclusion
In this paper, we conduct the first in-depth and comprehensive analysis to reveal the security vulnerabilities of backdoored PEFT-based adapters. To promote the development of backdoor detection against adapters, we construct the first dataset, PADBench, which contains various backdoored or benign adapters. Meanwhile, we propose the first backdoor detection method, named $\mathsf{PEFTGuard}$. $\mathsf{PEFTGuard}$ does not require any additional information during the detection. Instead, it only requires access to the parameters of the PEFT adapter. $\mathsf{PEFTGuard}$ achieves state-of-the-art performance in detecting various types of adapters, which are generated from different datasets, backdoor attack algorithms, PEFT methods, and various base models (including different model parameter sizes, both base and fine-tuned models, diverse model structures, and different attention mechanisms). In addition, $\mathsf{PEFTGuard}$ demonstrates a zero-shot transferability across different PEFT fine-tuning methods, adapter ranks, and backdoor attacks. Furthermore, we show that $\mathsf{PEFTGuard}$ is robust against different adaptive attacks. Overall, we hope that the proposed PADBench and $\mathsf{PEFTGuard}$ will play a key role in advancing the security governance of adapters in the open-source platforms.
Appendix A Ablation Study about Zero-shot Transferability on Unknown Attacks.
Table 14: Ablation Study Performance of Zero-Shot Transferability Analysis. (Notation: Sent = InsertSent, Word = RIPPLES, Syn = Syntactic, Sty = StyleBkd)
Training Attack Fusion (models) CL Detection Acc (%) AUC
Sent+Sty ✓, $3$ $59.00$ $0.973$
Sent+Word ✓, $3$ $50.00$ $0.815$
Word+Sty ✓, $3$ $57.00$ $0.971$
Sent+Word+Sty - $68.00$ $0.972$
Sent+Word+Sty ✓, $3$ $74.00$ $0.936$
Sent+Word+Sty ✓, $2$ $85.00$ $0.988$
Sent+Word+Sty ✓, $3$ $91.00$ $0.993$
In Section 6.4, we employ Contrastive Learning (CL), combined with model fusion via parameter averaging, enabling models trained on three different attack datasets to generalize to another unknown attack dataset. To investigate the effectiveness of each component, we conduct ablation studies, specifically targeting the unknown attack as Syntactic. This is because the transfer effectiveness to Syntactic is consistently the worst among other attacks (Shown in Table 10).
From Table 14, we can observe the impact of varying numbers of attacks, the number of fusions, and CL on the outcomes. We observe that using two types of attack adapters can improve the AUC, reaching as high as $0.973$ but with a low DA $59\%$. Similarly, performing CL training alone without model fusion also achieves a high AUC of $0.973$, but the DA only reaches $68\%$. Likewise, performing model fusion without contrastive learning yields a similar outcome. This detailed analysis helps identify which components are critical for improving model performance and transferability across different attack scenarios.
Appendix B Related Work of Backdoor Mitigation Method.
The goal of backdoor mitigation methods is to directly eliminate backdoors from models, and these methods can be used in conjunction with detection techniques to first identify and then remove the backdoors. In this paper, we focus on methods to eliminate backdoors through training techniques [70, 54, 51, 75].
Experiments by Yao et al. [70] and Sha et al. [54] both indicate that fine-tuning backdoored models on a clean subset of training samples can mitigate the backdoors. Rafailov et al. [51] propose Direct Preference Optimization (DPO), an optimization method specifically designed for LLMs. This method utilizes the mapping relationship between reward functions and optimal policies, demonstrating that this constrained reward maximization problem can be accurately optimized through single-stage policy training. By setting texts that contain backdoor triggers but have normal answers as preferred texts, the DPO method can effectively eliminate backdoor influences in the model. Zhang et al. [75] propose the Fine-mixing method, which considers the clean pre-trained model weights before fine-tuning on clean data and mixes the backdoored weights with clean pre-trained weights. In addition, they utilize Embedding Purification (E-PUR) to detect and eliminate potential backdoor techniques within embeddings.
Table 15: Detailed parameter settings for the C&W attack.
Parameter Set Settings
$P_{1}$ $c=1\times 10^{-4}$, $\kappa=0$, $\text{iter}=20$, $\text{lr}=1\times 10^{-5}$
$P_{2}$ $c=5\times 10^{-3}$, $\kappa=0$, $\text{iter}=20$, $\text{lr}=5\times 10^{-4}$
$P_{3}$ $c=0.1$, $\kappa=0$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{4}$ $c=0.1$, $\kappa=5$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{5}$ $c=0.5$, $\kappa=10$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
Figure 7: Architecture of $\mathsf{PEFTGuard}_{\rm cls}$.
Figure 8: Alternative Architectures of $\mathsf{PEFTGuard}$, representing two examples among many potential configurations.
Table 16: Details of our PADBench Datasets. “-” indicates empty because the instruction-following datasets can’t be evaluated for clean accuracy. “*” indicates that in IMDB and AG News datasets, the benign adapters remain the same across different attacks in the same training dataset.(Unless specified otherwise, the target projection matrices in the PEFT method are, by default, applied to the query and value matrices.)
Base Model Dataset Attack Method PEFT Method Rank Number Clean Accuracy ASR
Benign/Backdoored Benign Backdoored
Llama-2-7B SQuAD InsertSent LoRA $256$ $/250$ 250 $0.647$ $0.661$ $0.997$
Llama-2-7B toxic-backdoors-alpaca Word LoRA $256$ $/250$ 250 - - $0.964$
Llama-2-7B IMDB RIPPLES LoRA $256$ $250250^{*}/$ 250 $0.970$ $0.973$ $0.943$
InsertSent LoRA $256$ $250250^{*}/$ 250 $0.970$ $0.970$ $0.998$
Syntactic LoRA $256$ $250250^{*}/$ 250 $0.970$ $0.969$ $0.987$
StyleBkd LoRA $256$ $250250^{*}/$ 250 $0.970$ $0.960$ $0.949$
Llama-2-7B AG News RIPPLES LoRA $256$ $250250^{*}/$ 250 $0.940$ $0.940$ $0.948$
InsertSent LoRA $256$ $250250^{*}/$ 250 $0.940$ $0.938$ $0.969$
Syntactic LoRA $256$ $250250^{*}/$ 250 $0.940$ $0.940$ $0.986$
StyleBkd LoRA $256$ $250250^{*}/$ 250 $0.940$ $0.943$ $0.927$
Llama-2-7B toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.926$
Sentence LoRA (q) $512$ $/250$ 250 - - $0.938$
Sentence LoRA (k) $512$ $/250$ 250 - - $0.936$
Sentence LoRA (v) $512$ $/250$ 250 - - $0.942$
Sentence LoRA (q,k) $256$ $/250$ 250 - - $0.934$
Sentence LoRA (q,k,v,o) 128 $/250$ 250 - - $0.979$
Sentence QLoRA $256$ $/250$ 250 - - $0.796$
Sentence DoRA $256$ $/250$ 250 - - $0.787$
Sentence LoRA+ $8$ $/250$ 250 - - $0.644$
Sentence AdaLoRA $8$ $/250$ 250 - - $0.112$
Sentence LoRA $8$ $/50$ 50 - - $0.585$
Sentence LoRA $16$ $/50$ 50 - - $0.707$
Sentence LoRA $32$ $/50$ 50 - - $0.739$
Sentence LoRA $64$ $/50$ 50 - - $0.816$
Sentence LoRA $128$ $/50$ 50 - - $0.734$
Sentence LoRA $512$ $/50$ 50 - - $0.846$
Sentence LoRA $1024$ $/50$ 50 - - $0.833$
Sentence LoRA $2048$ $/50$ 50 - - $0.814$
Llama-2-13B toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.835$
Llama-3-8B toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.843$
Qwen1.5-7B-Chat toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.677$
ChatGLM-6B-v2 toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.641$
flan-t5-xl toxic-backdoors-hard Sentence LoRA $256$ $/250$ 250 - - $0.479$
Roberta-base IMDB InsertSent LoRA $256$ $/250$ 250 $0.955$ $0.950$ $1.000$
Qwen2-vl-2B VQAv2 InsertSent LoRA $16$ $/250$ 250 $0.735$ $0.738$ $0.649$
ViT-base CIFAR-10 BadNet LoRA $16$ $/250$ 250 $0.985$ $0.984$ $0.921$
Table 17: Overview of Backdoor Injection Datasets and Tasks
Type Dataset Task
Task-Specific SQuAD [53] Question Answering (QA)
AG News [74] Topic Classification
IMDB Movie [41] Sentiment Classification (SC)
CIFAR-10 [31] Image Classification
VQAv2 [18] Visual Question Answering
Instruction-Following (IF) toxic-backdoors-hard [16] Generation
toxic-backdoors-alpaca [15] Generation